# Untangling Lariats: Subgradient Following of Variationally Penalized Objectives

### v1

@@ 4 Variationally Penalized Multivariate Regression
We now extend the subgradient following framework to multivariate settings. For concreteness we examine objectives for instance $i$ of the form,
$h_{i}({\mathbf{x}})\,=\,\frac{1}{2}\big{\|}{\mathbf{x}}-\mathbf{y}_{i}\big{\|}^{2}_{{Q}_{i}}\,=\,\frac{1}{2}({\mathbf{x}}-\mathbf{y}_{i})^{\top}\mathsf{Q}_{i}({\mathbf{x}}-\mathbf{y}_{i})~{}~{}{\mathbf{x}},\mathbf{y}_{i}\in\mathbb{R}^{n}~{}.$ (10)
@@
We impose the condition that the quadratic form is strictly positive definite, $\mathsf{Q}_{i}\succ 0$. The multivariate is intrinsically more complex than the scalar case due to the penalty entanglement of ${\mathbf{x}}_{i}$ with itself and with ${\mathbf{x}}_{i+1}$. To mitigate this difficulty we examine first the squared Euclidean distance between ${\mathbf{x}}_{i}$ and ${\mathbf{x}}_{i+1}$,
$g_{i}(\boldsymbol{\delta})\,=\,\frac{1}{2}\lVert{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\rVert^{2}\,=\,\frac{1}{2}\lVert\boldsymbol{\delta}\rVert^{2}~{}.$
@@
This setting generalizes squared fused penalty for the scalar case for which $Q_{i}=[q_{i}\in\mathbb{R}]$ and thus can be absorbed into $\lambda_{i}$. For this setting we have,
$\nabla h_{i}({\mathbf{x}})=\mathsf{Q}_{i}({\mathbf{x}}-\mathbf{y}_{i})~{}\mbox{and}~{}\nabla g_{i}^{*}(\boldsymbol{\delta})=\boldsymbol{\delta}~{}.$
@@
The derivation of subgradient following for this setting also serves as a building block in the more complex multivariate settings presented in the sequel.
The Fenchel dual in the multivariate case where $f:\mathbb{R}^{n}\to\mathbb{R}$ is defined as,
$f^{*}(\boldsymbol{\alpha})=\sup_{{\mathbf{x}}\in\mathrm{dom}(f)}\big{\{}\boldsymbol{\alpha}^{\top}{\mathbf{x}}-f({\mathbf{x}})\big{\}}~{}.$
@@
Since both $h_{i}(\cdot)$ and $g_{i}(\cdot)$ defined above are strictly convex, as we require $\mathsf{Q}_{i}\succ 0$ the Fenchel dual for all $i\in[n]$ are unique and satisfy for all feasible dual pairs $({\mathbf{x}},\boldsymbol{\alpha})$ the equation $\boldsymbol{\alpha}=\nabla\!f({\mathbf{x}})$ typically referred to as the link function or the mirror map. We use this property in the ensuing derivation.
@@
Since the gradients’ components are of unary form and linear the subgradient sets can be written succinctly as,
$S_{i}=\left\{({\mathbf{x}},\boldsymbol{\mathsf{a}}):\mathsf{A}_{i}{\mathbf{x}}+\mathsf{B}_{i}\boldsymbol{\mathsf{a}}+{\mathbf{c}}_{i}=\mathbf{0}\right\}~{}.$
Since the mirror map here is bijective the admissible sets of (6) and (6) can be written as,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\boldsymbol{\mathsf{a}}_{i}=\nabla\!\tilde{f}_{i+1}(\boldsymbol{\mathsf{x}}_{i+1})\big{\}}$
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):\boldsymbol{\mathsf{a}}_{i}=\nabla\!f_{i}(\boldsymbol{\mathsf{x}}_{i})\big{\}}~{}.$
@@
We therefore can easily unravel the recursive form for $\mathsf{A}_{i-1}$, $\mathsf{B}_{i-1}$, ${\mathbf{c}}_{i-1}$ from $\mathsf{A}_{i}$, $\mathsf{B}_{i}$, ${\mathbf{c}}_{i}$,
@@
$S_{i-\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):\boldsymbol{\mathsf{a}}-\boldsymbol{\mathsf{a}}^{\prime}=\mathsf{Q}_{i}(\boldsymbol{\mathsf{x}}-\mathbf{y}_{i});\;(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}^{\prime})\in S_{i}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):\mathsf{A}_{i}\boldsymbol{\mathsf{x}}+\mathsf{B}_{i}\left(\boldsymbol{\mathsf{a}}-\mathsf{Q}_{i}(\boldsymbol{\mathsf{x}}-\mathbf{y}_{i})\right)+{\mathbf{c}}_{i}=\mathbf{0}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):(\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i})\boldsymbol{\mathsf{x}}+\mathsf{B}_{i}\boldsymbol{\mathsf{a}}+({\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i})=\mathbf{0}\big{\}}~{},$ $\vskip 3.0pt plus 1.0pt minus 1.0ptS_{i-1}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):\boldsymbol{\mathsf{x}}-\boldsymbol{\mathsf{x}}^{\prime}=\boldsymbol{\mathsf{a}};\;(\boldsymbol{\mathsf{x}}^{\prime},\boldsymbol{\mathsf{a}})\in S_{i-\frac{1}{2}}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):(\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i})(\boldsymbol{\mathsf{x}}-\boldsymbol{\mathsf{a}})+\mathsf{B}_{i}\boldsymbol{\mathsf{a}}+({\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i})=\mathbf{0}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):(\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i})\boldsymbol{\mathsf{x}}+(\mathsf{B}_{i}(\mathsf{I}+\mathsf{Q}_{i})-\mathsf{A}_{i})\boldsymbol{\mathsf{a}}+({\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i})=\mathbf{0}\big{\}}~{}.$
We therefore get that,
@@
${\mathbf{c}}_{i-1}$ $=$ ${\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i}~{}.$
@@
$(\mathsf{A}_{n},\mathsf{B}_{n},{\mathbf{c}}_{n})=(\mathsf{O},\,\mathsf{I},\,\mathbf{0})~{}\mbox{ and }~{}(\mathsf{A}_{1}-\mathsf{B}_{1}\mathsf{Q}_{1}){\mathbf{x}}_{1}+({\mathbf{c}}_{1}+\mathsf{B}_{1}\mathsf{Q}_{1}\mathbf{y}_{1})=0~{}.$
Upon obtaining ${\mathbf{x}}_{1}$ along with $\boldsymbol{\alpha}_{0}\equiv\mathbf{0}$, we construct the trajectory of the solution by following a backward recursion,
@@
${\mathbf{x}}_{i+1}$ $=$ ${\mathbf{x}}_{i}-\boldsymbol{\alpha}_{i}~{}.$
@@ 5 Sparse Variationally Penalized Multivariate Regression
We use ${\mathbf{x}}_{1:n}$ henceforth to denote the sequence of vectors ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$. While the algorithm of Sec. 4 entertains a concise form it does not yield variational sparsity. To do so we can readily use the $1$-norm for the variational penalty. This would yield though element-wise fusion whereas full variational sparsity entails that for some indices $i\in[n]$, ${\mathbf{x}}_{i}={\mathbf{x}}_{i+1}$. In this section we use the algorithm from as a building block while replace the norm squared of ${\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}$ with either the $2$-norm of the difference vector or its infinity norm. For brevity, we assume throughout this section that $\forall i,\lambda_{i}=1$ and are thus omitted.
@@ 5.1 Multivariate Regression with $\ell_{2}$-Variational Penalty
We first replace each of the norm-squared variational penalties with a norm, thus the goal is to find the minimizer of,
@@
Alas, this formulation does not entail a closed form subgradient following procedure. The approach that we take here is to construct a relaxation that yields the follow convergent algorithm.
Algorithm 1 Surrogate Multivariate Nariat
@@
$g_{\epsilon}(\boldsymbol{\delta})=\begin{cases}\frac{\|\boldsymbol{\delta}\|^{2}}{2\epsilon}&\|\boldsymbol{\delta}\|\leq\epsilon\\ \|\boldsymbol{\delta}\|-\frac{\epsilon}{2}&\|\boldsymbol{\delta}\|>\epsilon\end{cases}~{}.$
Since for any $\boldsymbol{\delta}$ it holds that $\|\boldsymbol{\delta}\|-\frac{\epsilon}{2}\leq g_{\epsilon}(\boldsymbol{\delta})\leq\|\boldsymbol{\delta}\|$, the solution of the relaxed problem Eq. (12) is at most $\frac{(n-1)\epsilon}{2}$ suboptimal with respect to the original problem defined in Eq. (11). Let us introduce a further auxiliary function,
$\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta}^{\prime})=\begin{cases}\frac{\|\boldsymbol{\delta}\|^{2}}{2\epsilon}&\|\boldsymbol{\delta}^{\prime}\|\leq\epsilon\\ \frac{\|\boldsymbol{\delta}\|^{2}}{2\|\boldsymbol{\delta}^{\prime}\|}+\frac{\|\boldsymbol{\delta}^{\prime}\|-\epsilon}{2}&\|\boldsymbol{\delta}^{\prime}\|>\epsilon\end{cases}~{}.$
The function $\hat{g}_{\epsilon}$ is a quadratic upper bound of $g_{\epsilon}(\boldsymbol{\delta})$ expanded at $\boldsymbol{\delta}^{\prime}$,
$\forall\boldsymbol{\delta},\boldsymbol{\delta}^{\prime}:~{}\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta}^{\prime})\geq g_{\epsilon}(\boldsymbol{\delta})~{}\mbox{ and }~{}\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta})=g_{\epsilon}(\boldsymbol{\delta})~{}.$
By carefully examining each cases we have the following,
@@
$\leq$ $g_{\epsilon}(\boldsymbol{\delta})+\frac{1}{2\,\max\big{(}\|\boldsymbol{\delta}^{\prime}\|,\epsilon\big{)}}\|\boldsymbol{\delta}-\boldsymbol{\delta}^{\prime}\|^{2}~{}.$
Hence, Algorithm 1 is a proximal method where we iteratively apply the multivariate subgradient following from previous section to obtain an exact solution for each adaptively constructed upper bound. We follow a similar skeletal proof to show the convergence.
We denote by $\eta^{t}=\min_{i}\rho_{i}^{t}\geq\epsilon$ and the optimum of the relaxed problem as $\hat{\phi}^{*}=\hat{\phi}({\mathbf{x}}_{1:n}^{*})$. We next prove a sequence of bounds as the means to show convergence starting with employing the definition of $g_{\epsilon}$ and $\hat{g}_{\epsilon}$,
@@
$\leq$ $\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})+\sum_{i=1}^{n}\frac{2}{\eta^{t}}\|{\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\|^{2}~{}.$
Next we introduce a dummy sequence ${\mathbf{z}}_{1:n}(\mu)=(1-\mu)\,{\mathbf{x}}_{1:n}^{t}+\mu\,{\mathbf{x}}_{1:n}^{*}$ and upper bound the last term by a minimizer constrained to the segment from ${\mathbf{x}}_{1:n}^{t}$ to ${\mathbf{x}}_{1:n}^{*}$ and then use the convexity of the objective,
@@
$\leq$ $\min_{\mu\in[0,1]}(1-\mu)\,\hat{\phi}({\mathbf{x}}_{1:n}^{t})+\mu\,\hat{\phi}({\mathbf{x}}_{1:n}^{*})+\sum_{i=1}^{n}\Big{(}\frac{2\mu^{2}}{\eta^{t}}-\frac{\mu(1-\mu)}{2}\Big{)}\|{\mathbf{x}}_{i}^{t}-{\mathbf{x}}_{i}^{*}\|^{2}~{}.$
Last, we replace the minimizer w.r.t. $\mu$ with a specific choice $\mu=\frac{\eta^{t}}{4+\eta^{t}}$ and get,
$\hat{\phi}({\mathbf{x}}_{1:n}^{t+1})$$\leq$ $\frac{4}{4+\eta^{t}}\hat{\phi}({\mathbf{x}}_{1:n}^{t})+\frac{\eta^{t}}{4+\eta^{t}}\hat{\phi}({\mathbf{x}}_{1:n}^{*})~{}.$
@@
$\hat{\phi}({\mathbf{x}}_{1:n}^{t+1})-\hat{\phi}^{*}\leq\frac{4}{4+\eta^{t}}\big{(}\hat{\phi}({\mathbf{x}}_{1:n}^{t})-\hat{\phi}^{*}\big{)}\leq\frac{4}{4+\epsilon}\big{(}\hat{\phi}({\mathbf{x}}_{1:n}^{t})-\hat{\phi}^{*}\big{)}~{}.$
Thus,we get that the sequence $\hat{\phi}({\mathbf{x}}_{1:n}^{t})$ convergence linearly to $\hat{\phi}^{*}$.

### v2

@@ 4 Variationally Penalized Multivariate Regression
We next extend the subgradient following framework to multivariate settings. For concreteness we examine objectives for instance $i$ of the form,
$h_{i}({\mathbf{x}})\,=\,\frac{1}{2}\big{\|}{\mathbf{x}}-\mathbf{y}_{i}\big{\|}^{2}_{{Q}_{i}}\,=\,\frac{1}{2}({\mathbf{x}}-\mathbf{y}_{i})^{\top}\mathsf{Q}_{i}({\mathbf{x}}-\mathbf{y}_{i})~{}~{}{\mathbf{x}},\mathbf{y}_{i}\in\mathbb{R}^{n}~{}~{}.$ (10)
@@
We assume w.l.o.g that the quadratic form is strictly positive definite, $\mathsf{Q}_{i}\succ 0$. The multivariate setting is intrinsically more complex than the scalar case due to the penalty entanglement of ${\mathbf{x}}_{i}$ with itself and with ${\mathbf{x}}_{i+1}$. To mitigate this difficulty we first examine the squared Euclidean distance between ${\mathbf{x}}_{i}$ and ${\mathbf{x}}_{i+1}$,
$g_{i}(\boldsymbol{\delta})\,=\,\frac{1}{2}\lVert{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\rVert^{2}\,=\,\frac{1}{2}\lVert\boldsymbol{\delta}\rVert^{2}~{}~{}.$
@@
This setting generalizes the squared fused penalty of the scalar case for which $Q_{i}=[q_{i}\in\mathbb{R}]$ and thus can be absorbed into $\lambda_{i}$. With these assumptions we get,
$\nabla h_{i}({\mathbf{x}})=\mathsf{Q}_{i}({\mathbf{x}}-\mathbf{y}_{i})~{}\mbox{and}~{}\nabla g_{i}^{*}(\boldsymbol{\delta})=\boldsymbol{\delta}~{}~{}.$
@@
The derivation of the subgradient following for this setting would also serve us as a building block in more complex multivariate problems described and analyzed in the sequel.
The Fenchel dual for the multivariate case where $f:\mathbb{R}^{n}\to\mathbb{R}$ is defined as,
$f^{*}(\boldsymbol{\alpha})=\sup_{{\mathbf{x}}\in\mathrm{dom}(f)}\big{\{}\boldsymbol{\alpha}^{\top}{\mathbf{x}}-f({\mathbf{x}})\big{\}}~{}~{}.$
@@
Both $h_{i}(\cdot)$ and $g_{i}^{*}(\cdot)$ defined above are strictly convex, as we require $\mathsf{Q}_{i}\succ 0$, thus the Fenchel duals for all $i\in[n]$ have a unique maximizer that forms dual feasible pairs $({\mathbf{x}},\boldsymbol{\alpha})$. The equation $\boldsymbol{\alpha}=\nabla\!f({\mathbf{x}})$ is typically referred to as the link function or mirror map. We use this property in the ensuing derivation.
@@
Since the components of the gradients are of unary form and linear, the subgradient sets can be written succinctly as,
$S_{i}=\left\{({\mathbf{x}},\boldsymbol{\mathsf{a}}):\mathsf{A}_{i}{\mathbf{x}}+\mathsf{B}_{i}\boldsymbol{\mathsf{a}}+{\mathbf{c}}_{i}=\mathbf{0}\right\}~{}~{}.$
Furthermore, as the mirror map here is bijective, the admissible sets of Eq. (5) and Eq. (6) can be written as,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\boldsymbol{\mathsf{a}}_{i}=\nabla\!f_{i+\frac{1}{2}}(\boldsymbol{\mathsf{x}}_{i+1})\big{\}}$
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):\boldsymbol{\mathsf{a}}_{i}=\nabla\!f_{i}(\boldsymbol{\mathsf{x}}_{i})\big{\}}~{}~{}.$
@@
We therefore effortlessly unravel the recursive form for $\mathsf{A}_{i-1}$, $\mathsf{B}_{i-1}$, ${\mathbf{c}}_{i-1}$ from $\mathsf{A}_{i}$, $\mathsf{B}_{i}$, ${\mathbf{c}}_{i}$,
@@
$S_{i-\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):\boldsymbol{\mathsf{a}}-\boldsymbol{\mathsf{a}}^{\prime}=\mathsf{Q}_{i}(\boldsymbol{\mathsf{x}}-\mathbf{y}_{i});\;(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}^{\prime})\in S_{i}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):\mathsf{A}_{i}\boldsymbol{\mathsf{x}}+\mathsf{B}_{i}\left(\boldsymbol{\mathsf{a}}-\mathsf{Q}_{i}(\boldsymbol{\mathsf{x}}-\mathbf{y}_{i})\right)+{\mathbf{c}}_{i}=\mathbf{0}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):(\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i})\boldsymbol{\mathsf{x}}+\mathsf{B}_{i}\boldsymbol{\mathsf{a}}+({\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i})=\mathbf{0}\big{\}}~{}~{},$ $\vskip 3.0pt plus 1.0pt minus 1.0ptS_{i-1}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):\boldsymbol{\mathsf{x}}-\boldsymbol{\mathsf{x}}^{\prime}=\boldsymbol{\mathsf{a}};\;(\boldsymbol{\mathsf{x}}^{\prime},\boldsymbol{\mathsf{a}})\in S_{i-\frac{1}{2}}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):(\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i})(\boldsymbol{\mathsf{x}}-\boldsymbol{\mathsf{a}})+\mathsf{B}_{i}\boldsymbol{\mathsf{a}}+({\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i})=\mathbf{0}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):(\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i})\boldsymbol{\mathsf{x}}+(\mathsf{B}_{i}(\mathsf{I}+\mathsf{Q}_{i})-\mathsf{A}_{i})\boldsymbol{\mathsf{a}}+({\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i})=\mathbf{0}\big{\}}~{}~{}.$
In summary, we get that,
@@
${\mathbf{c}}_{i-1}$ $=$ ${\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i}~{}~{}.$
@@
$(\mathsf{A}_{n},\mathsf{B}_{n},{\mathbf{c}}_{n})=(\mathsf{O},\,\mathsf{I},\,\mathbf{0})~{}\mbox{ and }~{}(\mathsf{A}_{1}-\mathsf{B}_{1}\mathsf{Q}_{1}){\mathbf{x}}_{1}+({\mathbf{c}}_{1}+\mathsf{B}_{1}\mathsf{Q}_{1}\mathbf{y}_{1})=0~{}~{}.$
Upon obtaining ${\mathbf{x}}_{1}$ along with $\boldsymbol{\alpha}_{0}\equiv\mathbf{0}$, we construct the trajectory of the solution using the backward recursion,
@@
${\mathbf{x}}_{i+1}$ $=$ ${\mathbf{x}}_{i}-\boldsymbol{\alpha}_{i}~{}~{}.$
@@ 5 Sparse Variationally Penalized Multivariate Regression
We henceforth use ${\mathbf{x}}_{1:n}$ to denote the sequence of vectors ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$. While the algorithm of Sec. 4 entertains a concise form, it does not yield variational sparsity. To do so we can readily use the $1$-norm as the variational penalty. This would yield though element-wise fusion whereas full variational sparsity entails that for some indices $i\in[n]$, ${\mathbf{x}}_{i}={\mathbf{x}}_{i+1}$. In this section we use the algorithm from the previous section as a building block while replacing the norm-squared penalty $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|^{2}$ with either the $2$-norm $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$ of the difference vector or its infinity norm $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}$. For brevity, we assume throughout this section that $\forall i,\lambda_{i}=1$ and are thus omitted.
@@ 5.1 Multivariate Regression with $\ell_{2}$-Variational Penalty
When we replace each of the norm-squared variational penalties with the $2$-norm, the goal is to find the minimizer of,
@@
Alas, this formulation does not entail a closed form subgradient following procedure. The approach that we take here is to construct a relaxation that yields the following convergent algorithm.
Algorithm 1 Surrogate for Multivariate Untangled Lariat
@@
$g_{\epsilon}(\boldsymbol{\delta})=\begin{cases}\frac{\|\boldsymbol{\delta}\|^{2}}{2\epsilon}&\|\boldsymbol{\delta}\|\leq\epsilon\\ \|\boldsymbol{\delta}\|-\frac{\epsilon}{2}&\|\boldsymbol{\delta}\|>\epsilon\end{cases}~{}~{}.$
Since for any $\boldsymbol{\delta}$ it holds that $\|\boldsymbol{\delta}\|-\frac{\epsilon}{2}\leq g_{\epsilon}(\boldsymbol{\delta})\leq\|\boldsymbol{\delta}\|$, the solution of the relaxed problem Eq. (12) is at most $\frac{(n-1)\epsilon}{2}$ suboptimal with respect to that of the original problem defined in Eq. (11). Let us introduce a further auxiliary function,
$\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta}^{\prime})=\begin{cases}\frac{\|\boldsymbol{\delta}\|^{2}}{2\epsilon}&\|\boldsymbol{\delta}^{\prime}\|\leq\epsilon\\ \frac{\|\boldsymbol{\delta}\|^{2}}{2\|\boldsymbol{\delta}^{\prime}\|}+\frac{\|\boldsymbol{\delta}^{\prime}\|-\epsilon}{2}&\|\boldsymbol{\delta}^{\prime}\|>\epsilon\end{cases}~{}~{}.$
The function $\hat{g}_{\epsilon}$ is a quadratic upper bound on $g_{\epsilon}(\boldsymbol{\delta})$ expanded at $\boldsymbol{\delta}^{\prime}$, thus $\forall\boldsymbol{\delta},\boldsymbol{\delta}^{\prime}:~{}\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta}^{\prime})\geq g_{\epsilon}(\boldsymbol{\delta})$ and $\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta})=g_{\epsilon}(\boldsymbol{\delta})$. By examining each case we get the following,
@@
$\leq$ $g_{\epsilon}(\boldsymbol{\delta})+\frac{1}{2\,\max\big{(}\|\boldsymbol{\delta}^{\prime}\|,\epsilon\big{)}}\|\boldsymbol{\delta}-\boldsymbol{\delta}^{\prime}\|^{2}~{}~{}.$
Hence, Algorithm 1 is a proximal method where we iteratively apply the multivariate subgradient-following of the previous section to obtain an exact solution for each adaptively constructed upper bound. We follow a similar skeletal proof to show the convergence.
We denote by $\eta^{t}=\min_{i}\rho_{i}^{t}\geq\epsilon$ and the optimum of the relaxed problem as $\hat{\phi}^{*}=\hat{\phi}({\mathbf{x}}_{1:n}^{*})$. We derive a series of bounds as the means to show convergence, starting with the definitions of $g_{\epsilon}$ and $\hat{g}_{\epsilon}$,
@@
$\leq$ $\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})+\sum_{i=1}^{n}\frac{2}{\eta^{t}}\|{\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\|^{2}~{}~{}.$
Next we introduce a auxiliary sequence ${\mathbf{z}}_{1:n}(\mu)=(1-\mu)\,{\mathbf{x}}_{1:n}^{t}+\mu\,{\mathbf{x}}_{1:n}^{*}$ and upper bound the last inequality using a minimizer constrained to the line segment from ${\mathbf{x}}_{1:n}^{t}$ to ${\mathbf{x}}_{1:n}^{*}$. We then use the convexity of the objective. These two steps amount to,
@@
$\leq$ $\min_{\mu\in[0,1]}(1-\mu)\,\hat{\phi}({\mathbf{x}}_{1:n}^{t})+\mu\,\hat{\phi}({\mathbf{x}}_{1:n}^{*})+\sum_{i=1}^{n}\Big{(}\frac{2\mu^{2}}{\eta^{t}}-\frac{\mu(1-\mu)}{2}\Big{)}\|{\mathbf{x}}_{i}^{t}-{\mathbf{x}}_{i}^{*}\|^{2}~{}~{}.$
Finally, we replace the minimizer w.r.t. $\mu$ with a specific choice $\mu=\frac{\eta^{t}}{4+\eta^{t}}$ and get,
$\hat{\phi}({\mathbf{x}}_{1:n}^{t+1})$$\leq$ $\frac{4}{4+\eta^{t}}\hat{\phi}({\mathbf{x}}_{1:n}^{t})+\frac{\eta^{t}}{4+\eta^{t}}\hat{\phi}({\mathbf{x}}_{1:n}^{*})~{}~{}.$
@@
$\hat{\phi}({\mathbf{x}}_{1:n}^{t+1})-\hat{\phi}^{*}\leq\frac{4}{4+\eta^{t}}\big{(}\hat{\phi}({\mathbf{x}}_{1:n}^{t})-\hat{\phi}^{*}\big{)}\leq\frac{4}{4+\epsilon}\big{(}\hat{\phi}({\mathbf{x}}_{1:n}^{t})-\hat{\phi}^{*}\big{)}~{}~{}.$
In summary, we get that the sequence $\hat{\phi}({\mathbf{x}}_{1:n}^{t})$ converges linearly to $\hat{\phi}^{*}$.